The nitrogen dioxide content of air on a particular day was 0.150 ppm by mass.
The density of the air was 1.225kgm−3.
What was the mass of nitrogen dioxide in 1m3
of air on that day?
1
Expert's answer
2018-01-04T03:45:06-0500
Answer:
m = ρ · V = 1.225 · 1 = 1.225 kg
ppm = mg/kg
m (NO 2 ) = 0.150 · 1.225 = 0.184 mg